OpenStudy (anonymous):

I Give Medals!!!!!!!!!!!!!!!!!!!!1 Solve the equation for y. 8x – 9y = 11

OpenStudy (anonymous):

8x - 9y = 11 8x - 11 = 9y y = (8x-11)/9

OpenStudy (anonymous):

Which ordered pair in the form (a, b) is a solution of this equation? 7a – 5b = 28

OpenStudy (anonymous):

A. (–3, –2) B. (–2, –3) C. (4, 0) D. (0, 4)

OpenStudy (anonymous):

okay one second trying to figure this one out myself